Numerical Analysis I, Fall 2014 (http://www.math.nthu.edu.tw/˜ wangwc/)

Quiz 0

Sep 19, 2014.

All the answers can be found by ’for’ loop, ’if’ loop, ’while’ loop, ’and’, ’or’, etc.

1. Find the smallest number among sin n, n = 1, 2, ⋯ 100, and the corresponding n.

2. Find the largest N such that ∑_{j=1}^{N}(1/j) < 5 and the corresponding sum.

3. A point (i, j) is called a grid point if both i, j are integers. How many grid points are there in the set {(x, y)|(x−10)^{2}+(y−10)^{2} < 35}∪{(x, y)|(x−10)^{2}+(y−10)^{2}/4 < 50}?

Numerical Analysis I, Fall 2014 (http://www.math.nthu.edu.tw/˜ wangwc/)

Quiz 0

Sep 19, 2014.

All the answers can be found by ’for’ loop, ’if’ loop, ’while’ loop, ’and’, ’or’, etc.

1. Find the smallest number among sin n, n = 1, 2, ⋯ 100, and the corresponding n.

2. Find the largest N such that ∑_{j=1}^{N}(1/j) < 5 and the corresponding sum.

3. A point (i, j) is called a grid point if both i, j are integers. How many grid points are there in the set {(x, y)|(x−10)^{2}+(y−10)^{2} < 35}∪{(x, y)|(x−10)^{2}+(y−10)^{2}/4 < 50}?